# How Transformers Implement Induction Heads: Approximation and Optimization Analysis

### v2

@@ A.3.2 Proper Orthogonal Decomposition
Recently, Jiang and Li [2023] also used POD to study the approximation rate of single-layer single-head Transformer for the targets with nonlinear temporal kernels.
Given that two-layer FFNs can efficiently approximate Barron functions [Ma et al., 2020], which is dense in $L^{2}([0,1]^{d})$ [Siegel and Xu, 2020], we introduce the following technical definition regarding the well-behavior POD, which is used for our theoretical analysis.
Definition A.7 (Well-behaved POD).
@@
A.3.3 Proof of Theorem 4.4
$\textnormal{{GIH}}_{n}(X_{L})=\sum_{s=n}^{L-1}{\rm softmax}\Big{(}g\big{(}[x_{L},\cdots,x_{L-n+2}];[x_{s-1},\cdots,x_{s-n+1}]\big{)}\Big{)}x_{s},$ (16)
Theorem A.8 (Restatement of Theorem 4.4).
Let $\textnormal{{GIH}}_{n}$ satisfy Eq. (16). Suppose the similarity function $g$ is $\alpha$-well-behaved (see Definition A.7). Then, for any $q>0$, there exist constants $A_{g,q,n},B_{g,\alpha}>0$ and a two-layer $H$-head transformer ${\textnormal{{TF}}}(\cdot)$ with FFN of width $M$, such that the following approximation rate holds:
$\left|\!\left|\!\left|\textnormal{{GIH}}_{n}-{\textnormal{{TF}}}\right|\!\right|\!\right|_{L,2}\leq\frac{A_{g,q,n}}{H^{n}}+\frac{B_{g,\alpha}L^{1/(1+2\alpha)}}{M^{\alpha/(1+3\alpha)}}.$
@@
First, we set an constant $K\in\mathbb{N}^{+}$, and we will optimize it finally. We choose the embedding dimension $D=nd+2(n-1)K$, and the flowchart of the theorem proof is as follows:
$\sum_{s=n}^{L-1}{\rm softmax}\Big{(}\hat{g}_{K}$ $\big{(}\hat{X}_{L-n+2:L},\hat{X}_{s-n+1:s-1}\big{)}\Big{)}x_{s}$
@@
$,\displaystyle\Big{(}x_{L}^{\top},\ldots,\hat{x}_{L-n+1},\hat{\phi}_{1}(\hat{X}_{L-n+2:L}),$$\ldots,\hat{\phi}_{K}(\hat{X}_{L-n+2:L}),$
$\hat{\psi}_{1}(\hat{X}_{L-n+1:L-1}),$ $\ldots,\hat{\psi}_{K}(\hat{X}_{L-n+1:L-1})\Big{)}^{\top}$
@@
$(x_{t}^{\top},$ ${0}^{\top})^{\top}$
Recalling Definition A.7, there exists constants $C_{g}^{\infty},C_{g}^{\mathcal{B}},C_{g}^{{\rm Lip}}>0$ such that:
@@
Step I: Error in 1-st Attn layer. This step is essentially the same as Step I in the proof of Theorem 4.3, so we write down the error of the first Attn layer directly:
$\epsilon_{{\textnormal{{SA}}}}^{(1)}\leq\frac{C_{q,n}}{H^{n}}.$
Moreover, due to $\left\|\hat{X}_{s-n+2:s}-{X}_{s-n+2:s}\right\|\leq\epsilon_{{\textnormal{{SA}}}}^{(1)}$, for all $s$, we have:
$\hat{X}_{s-n+2:s}\in[-2,2]^{D}.$
Step II: Error in 1-st FFN layer. The $1$-st FFN is used to approximate $\phi_{k},\psi_{k}\ (k=1,\ldots,K)$. Each function is approximated by a $2$-layer neural networks with $\frac{M}{2K}$ neurons defined on $\mathbb{R}^{D}$, and the FFNs are concatenated together ( refer to section 7.1 "Parallelization" in  Schmidt-Hieber et al. [2020] ) as $\text{FFN}^{(1)}$. We denote them as
@@
Then according to lemma A.6, such FFNs exist and satisfy the following properties hold for all $1\leq k\leq K$:
$\|\hat{\phi}_{k}-\phi_{k}\|_{L^{\infty}([-2,2]^{D})}\leq\tilde{\mathcal{O}}\left(\|\phi_{k}\|_{\mathcal{B}}\sqrt{\frac{K}{{M}}}\right)\leq\epsilon_{\textnormal{{FFN}}}^{\rm(1)},$
$\|\hat{\psi}_{k}-\psi_{k}\|_{L^{\infty}([-2,2]^{D})}\leq\tilde{\mathcal{O}}\left(\|\psi_{k}\|_{\mathcal{B}}\sqrt{\frac{K}{{M}}}\right)\leq\epsilon_{\textnormal{{FFN}}}^{\rm(1)},$
@@
$\epsilon_{\textnormal{{FFN}}}^{(1)}:=\cdot\tilde{\mathcal{O}}\left(C_{g}^{\mathcal{B}}\sqrt{\frac{K}{{M}}}\right).$
@@
$\left\|x_{L}^{(2)}-\sum_{s=n}^{L-1}{\rm softmax}\Big{(}g\big{(}X_{L-n+2:L},X_{s-n+1:s-1}\big{)}\Big{)}x_{s}\right\|_{\infty}$ (17)
$=$ $\left\|\sum_{s=n}^{L-1}{\rm softmax}\Big{(}\hat{g}_{K}\big{(}\hat{X}_{L-n+2:L},\hat{X}_{s-n+1:s-1}\big{)}\Big{)}x_{s}-\sum_{s=n}^{L-1}{\rm softmax}\Big{(}g\big{(}X_{L-n+2:L},X_{s-n+1:s-1}\big{)}\Big{)}x_{s}\right\|_{\infty}$
$\leq$ $\left\|\sum_{s=n}^{L-1}{\rm softmax}\Big{(}\hat{g}_{K}\big{(}\hat{X}_{L-n+2:L},\hat{X}_{s-n+1:s-1}\big{)}\Big{)}x_{s}-\sum_{s=n}^{L-1}{\rm softmax}\Big{(}g_{K}\big{(}\hat{X}_{L-n+2:L},\hat{X}_{s-n+1:s-1}\big{)}\Big{)}x_{s}\right\|_{\infty}$
$+\left\|\sum_{s=n}^{L-1}{\rm softmax}\Big{(}g_{K}\big{(}\hat{X}_{L-n+2:L},\hat{X}_{s-n+1:s-1}\big{)}\Big{)}x_{s}-\sum_{s=n}^{L-1}{\rm softmax}\Big{(}g_{K}\big{(}{X}_{L-n+2:L},{X}_{s-n+1:s-1}\big{)}\Big{)}x_{s}\right\|_{\infty}$
$+\left\|\sum_{s=n}^{L-1}{\rm softmax}\Big{(}g_{K}\big{(}\hat{X}_{L-n+2:L},\hat{X}_{s-n+1:s-1}\big{)}\Big{)}x_{s}-\sum_{s=n}^{L-1}{\rm softmax}\Big{(}g\big{(}\hat{X}_{L-n+2:L},\hat{X}_{s-n+1:s-1}\big{)}\Big{)}x_{s}\right\|_{\infty}$
@@
$\int_{X\in[0,1]^{d\times L}}\left\|x_{L}^{(2)}-\sum_{s=n}^{L-1}{\rm softmax}\Big{(}g\big{(}X_{L-n+2:L},X_{s-n+1:s-1}\big{)}\Big{)}x_{s}\right\|_{\infty}^{2}{\textnormal{d}}X$
@@
$\leq$ $\mathcal{O}\left(\frac{C_{g,q,n}}{H^{n}}\right)+\tilde{\mathcal{O}}\left(\frac{C_{g,\alpha}\sqrt{K}}{\sqrt{M}}\right)+\mathcal{O}\left(\frac{tC_{\alpha}}{K^{\alpha}}\right)$
@@
Notice that in RHS of (18), only $\tilde{\mathcal{O}}\left(\frac{C_{g,\alpha}\sqrt{K}}{\sqrt{M}}\right)$ and $\mathcal{O}\left(\frac{LC_{\alpha}}{K^{\alpha}}\right)$ depend on $K$.
@@
$\min_{K}:\frac{\alpha}{\frac{1}{2}+\alpha}\frac{C_{g,\alpha}\sqrt{K}}{\sqrt{M}}+\frac{\frac{1}{2}}{\frac{1}{2}+\alpha}\frac{LC_{\alpha}}{K^{\alpha}}$
$=$ $\min_{K}:\frac{\alpha}{\frac{1}{2}+\alpha}\left(\left(\frac{C_{g,\alpha}\sqrt{K}}{\sqrt{M}}\right)^{\frac{\alpha}{\frac{1}{2}+\alpha}}\right)^{\frac{\frac{1}{2}+\alpha}{\alpha}}+\frac{\frac{1}{2}}{\frac{1}{2}+\alpha}\left(\left(\frac{LC_{\alpha}}{K^{\alpha}}\right)^{\frac{\frac{1}{2}}{\frac{1}{2}+\alpha}}\right)^{2(\frac{1}{2}+\alpha)}$
$=$ $\frac{C_{g,\alpha}^{\prime}L^{1/(1+2\alpha)}}{M^{\alpha/(1+2\alpha)}}.$
@@
$\left|\!\left|\!\left|\textnormal{{GIH}}_{n}-{\textnormal{{TF}}}\right|\!\right|\!\right|_{L,2}\leq\mathcal{O}\left(\frac{C_{g,q,n}}{H^{n}}\right)+\left\{\tilde{\mathcal{O}}\left(\frac{C_{g,\alpha}\sqrt{K}}{\sqrt{M}}\right)+\mathcal{O}\left(\frac{LC_{\alpha}}{K^{\alpha}}\right)\right\}_{\min:K}$
$\leq$ $\mathcal{O}\left(\frac{C_{g,q,n}}{H^{n}}\right)+\tilde{\mathcal{O}}\left(\frac{C_{g,\alpha}^{\prime}L^{1/(1+2\alpha)}}{M^{\alpha/(1+2\alpha)}}\right)\leq\frac{A_{g,q,n}}{H^{n}}+\frac{B_{g,\alpha}L^{1/(1+2\alpha)}}{M^{\alpha/(1+3\alpha)}}.$
Appendix B Proofs of Optimization Dynamics: Training Stage I

### v3

@@ C.3.2 Proper Orthogonal Decomposition
Recently, Jiang and Li (2023) also used POD to study the approximation rate of single-layer single-head Transformer for the targets with nonlinear temporal kernels.
Given that two-layer FFNs can efficiently approximate Barron functions (Ma et al., 2020), which is dense in $L^{2}([0,1]^{d})$ (Siegel and Xu, 2020), we introduce the following technical definition regarding the well-behavior POD, which is used for our theoretical analysis.
Definition C.7 (Well-behaved POD).
@@
C.3.3 Proof of Theorem 3.4
$\textnormal{{GIH}}_{n}(X_{L})=(x_{s})_{s=n}^{L-1}\ {\rm softmax}\Big{(}\big{(}g\big{(}X_{L-n+2:L};X_{s-n+1:s-1}\big{)}\big{)}_{s=n}^{L-1}\Big{)}^{\top},$ (16)
Theorem C.8 (Restatement of Theorem 3.4).
Suppose the similarity function $g$ is $\alpha$-well-behaved (see Definition C.7). Then there exist two absolute constants $A,B>0$ (only depending on the properties of $g$) such that: for any $H,M\in\mathbb{N}^{+}$ and rate $q\in\mathbb{N}^{+}$, there exists a constant $C_{n,q}>0$ and a two-layer $H$-head transformer ${\textnormal{{TF}}}(\cdot)$ with FFNs of width $M$, such that
$\left|\!\left|\!\left|\textnormal{{GIH}}_{n}-{\textnormal{{TF}}}\right|\!\right|\!\right|_{L,2}\leq A\left(\frac{C_{n,q}}{H}\right)^{q}+B\frac{L^{1/(1+2\alpha)}}{M^{\alpha/(1+3\alpha)}},$
where $C_{n,q}=\mathcal{O}(nq^{2})$.
@@
First, we set an constant $K\in\mathbb{N}^{+}$, and we will optimize it finally. We choose the embedding dimension $D=nd+2(n-1)K$.
Additionally, in this proof, the following projection matrices are used:
$P_{i}:=\left(0_{d\times(i-1)d},I_{d\times d},0_{d\times(D-id)}\right)\in\mathbb{R}^{d\times D},\quad i\in[n].$
The proof sketch can be summarized as follows:
$(x_{s})_{s=n}^{L-1}\ {\rm softmax}\Big{(}\big{(}g\big{(}X_{L-n+2:L}$ $;X_{s-n+1:s-1}\big{)}\big{)}_{s=n}^{L-1}\Big{)}^{\top}$
@@
$\Big{(}x_{L}^{\top},\ldots,\hat{x}_{L-n+1},\hat{\phi}_{1}(\hat{X}_{L-n+2:L}),\ldots,$ $\hat{\phi}_{K}(\hat{X}_{L-n+2:L}),\hat{\psi}_{1}(\hat{X}_{L-n+1:L-1}),\ldots,\hat{\psi}_{K}(\hat{X}_{L-n+1:L-1})\Big{)}^{\top}$
@@
$(x_{L}^{\top},$ ${0}^{\top})^{\top}$
Recalling Definition C.7, there exists constants $C_{g}^{\infty},C_{g}^{\mathcal{B}},C_{g}^{{\rm Lip}}>0$ such that:
@@
Step I: Error in 1-st Attn layer. This step is essentially the same as Step I in the proof of Theorem 3.3, so we write down the estimate of the first Attn layer directly: there exists absolute constants $C_{1},C_{2}>0$ such that
$\varepsilon_{{\textnormal{{SA}}}}^{(1)}:=\sup_{s}\left\|x_{s}^{(1)}-X_{s-n+1:s}\right\|_{\infty}\leq C_{2}\left(\frac{C_{1}nq^{2}}{H}\right)^{q}.$
Similar to the proof of Theorem 3.3, we first focus on the case of large $H$:
$H\geq C_{1}C_{2}enq^{2},$
which ensures $\varepsilon_{{\textnormal{{SA}}}}^{(1)}\leq 1$. Therefore, $x_{s}^{(1)}\in[-2,2]^{D}$:
For simplicity, we use the following notations:
$\hat{X}_{s-n+1:s-1}:=W_{Q}=\begin{pmatrix}P_{2}{x}_{s}^{(1)}\\ \vdots\\ P_{n}{x}_{s}^{(1)}\end{pmatrix},\ \text{ for }n\leq s\leq L-1;\quad\quad\hat{X}_{L-n+2:L}:=\begin{pmatrix}P_{1}{x}_{L}^{(1)}\\ \vdots\\ P_{n-1}{x}_{L}^{(1)}.\end{pmatrix}.$
Step II: Error in 1-st FFN layer. The $1$-st FFN is used to approximate $\phi_{k},\psi_{k}\ (k=1,\ldots,K)$. Each function is approximated by a $2$-layer neural networks with $\frac{M}{2K}$ neurons defined on $\mathbb{R}^{D}$, and the FFNs are concatenated together ( refer to section 7.1 "Parallelization" in  Schmidt-Hieber et al. (2020) ) as $\text{FFN}^{(1)}$. We denote them as
@@
Then by lemma C.6, such FFNs exist and satisfy the following properties hold for all $1\leq k\leq K$:
$\|\hat{\phi}_{k}-\phi_{k}\|_{L^{\infty}([-2,2]^{D})}\leq{\mathcal{O}}\left(\|\phi_{k}\|_{\mathcal{B}}\sqrt{\frac{K\log M}{{M}}}\right)\leq\epsilon_{\textnormal{{FFN}}}^{\rm(1)},$
$\|\hat{\psi}_{k}-\psi_{k}\|_{L^{\infty}([-2,2]^{D})}\leq{\mathcal{O}}\left(\|\psi_{k}\|_{\mathcal{B}}\sqrt{\frac{K\log M}{M}}\right)\leq\epsilon_{\textnormal{{FFN}}}^{\rm(1)},$
@@
$\epsilon_{\textnormal{{FFN}}}^{(1)}:=\mathcal{O}\left(C_{g}^{\mathcal{B}}\sqrt{\frac{K\log M}{{M}}}\right),\quad C_{g}^{\mathcal{B}}=\max\{\left\|\phi_{k}\right\|_{\mathcal{B}},\left\|\psi_{k}\right\|_{\mathcal{B}}\}.$
@@
$\left\|x_{L}^{(2)}-(x_{s})_{s=n}^{L-1}\ {\rm softmax}\Big{(}\big{(}g\big{(}X_{L-n+2:L};X_{s-n+1:s-1}\big{)}\big{)}_{s=n}^{L-1}\Big{)}^{\top}\right\|_{\infty}$ (17)
$\leq$ $\left\|{\rm softmax}\Big{(}\big{(}\hat{g}_{K}\big{(}\hat{X}_{L-n+2:L};\hat{X}_{s-n+1:s-1}\big{)}\big{)}_{s=n}^{L-1}\Big{)}^{\top}-{\rm softmax}\Big{(}\big{(}g\big{(}X_{L-n+2:L};X_{s-n+1:s-1}\big{)}\big{)}_{s=n}^{L-1}\Big{)}^{\top}\right\|_{\infty}$
$\leq$ $\left\|{\rm softmax}\Big{(}\big{(}\hat{g}_{K}\big{(}\hat{X}_{L-n+2:L};\hat{X}_{s-n+1:s-1}\big{)}\big{)}_{s=n}^{L-1}\Big{)}^{\top}-{\rm softmax}\Big{(}\big{(}g_{K}\big{(}\hat{X}_{L-n+2:L};\hat{X}_{s-n+1:s-1}\big{)}\big{)}_{s=n}^{L-1}\Big{)}^{\top}\right\|_{\infty}$
$+\left\|{\rm softmax}\Big{(}\big{(}g_{K}\big{(}\hat{X}_{L-n+2:L};\hat{X}_{s-n+1:s-1}\big{)}\big{)}_{s=n}^{L-1}\Big{)}^{\top}-{\rm softmax}\Big{(}\big{(}g_{K}\big{(}X_{L-n+2:L};X_{s-n+1:s-1}\big{)}\big{)}_{s=n}^{L-1}\Big{)}^{\top}\right\|_{\infty}$
$+\left\|{\rm softmax}\Big{(}\big{(}g_{K}\big{(}X_{L-n+2:L};X_{s-n+1:s-1}\big{)}\big{)}_{s=n}^{L-1}\Big{)}^{\top}-{\rm softmax}\Big{(}\big{(}g\big{(}X_{L-n+2:L};X_{s-n+1:s-1}\big{)}\big{)}_{s=n}^{L-1}\Big{)}^{\top}\right\|_{\infty}$
@@
$\int_{X\in[0,1]^{d\times L}}\left\|x_{L}^{(2)}-(x_{s})_{s=n}^{L-1}\ {\rm softmax}\Big{(}\big{(}g\big{(}X_{L-n+2:L};X_{s-n+1:s-1}\big{)}\big{)}_{s=n}^{L-1}\Big{)}^{\top}\right\|_{\infty}^{2}{\textnormal{d}}X$
@@
$\leq$ $\mathcal{O}\left(A_{g,\alpha}C_{2}\left(\frac{C_{1}nq^{2}}{H}\right)^{q}\right)+\mathcal{O}\left(\frac{B_{g,\alpha}\sqrt{K\log M}}{\sqrt{M}}\right)+\mathcal{O}\left(\frac{LC_{\alpha}}{K^{\alpha}}\right),$
where
$A_{g,\alpha}:=C_{g}^{\infty}C_{g}^{{\rm Lip}}C_{\alpha},\quad B_{g,\alpha}:=(2C_{g}^{\infty}+1)C_{\alpha}.$
@@
Notice that in RHS of (18), only ${\mathcal{O}}\left(\frac{C_{g,\alpha}\sqrt{K\log M}}{\sqrt{M}}\right)$ and $\mathcal{O}\left(\frac{LC_{\alpha}}{K^{\alpha}}\right)$ depend on $K$.
@@
$\min_{K}:\frac{\alpha}{\frac{1}{2}+\alpha}\frac{B_{g,\alpha}\sqrt{K\log M}}{\sqrt{M}}+\frac{\frac{1}{2}}{\frac{1}{2}+\alpha}\frac{LC_{\alpha}}{K^{\alpha}}$
$=$ $\min_{K}:\frac{\alpha}{\frac{1}{2}+\alpha}\left(\left(\frac{B_{g,\alpha}\sqrt{K\log M}}{\sqrt{M}}\right)^{\frac{\alpha}{\frac{1}{2}+\alpha}}\right)^{\frac{\frac{1}{2}+\alpha}{\alpha}}+\frac{\frac{1}{2}}{\frac{1}{2}+\alpha}\left(\left(\frac{LC_{\alpha}}{K^{\alpha}}\right)^{\frac{\frac{1}{2}}{\frac{1}{2}+\alpha}}\right)^{2(\frac{1}{2}+\alpha)}$
$=$ $\frac{B_{g,\alpha}^{\prime}L^{1/(1+2\alpha)}}{(M/\log M)^{\alpha/(1+2\alpha)}},$
where $B_{g,\alpha}^{\prime}$ only depends on the properties of $g$ and $\alpha$.
@@
$\left|\!\left|\!\left|\textnormal{{GIH}}_{n}-{\textnormal{{TF}}}\right|\!\right|\!\right|_{L,2}\leq\mathcal{O}\left(A_{g,q}C_{2}\left(\frac{C_{1}nq^{2}}{H}\right)^{q}\right)+\left\{{\mathcal{O}}\left(\frac{B_{g,\alpha}\sqrt{K\log M}}{\sqrt{M}}\right)+\mathcal{O}\left(\frac{LC_{\alpha}}{K^{\alpha}}\right)\right\}_{\min:K}$
$\leq$ $\mathcal{O}\left(A_{g,q}\left(\frac{C_{1}nq^{2}}{H}\right)^{q}\right)+{\mathcal{O}}\left(\frac{B_{g,\alpha}^{\prime}L^{1/(1+2\alpha)}}{(M/\log M)^{\alpha/(1+2\alpha)}}\right)$
$\leq$ $A_{g,q}^{\prime}\left(\frac{C_{1}nq^{2}}{H}\right)^{q}+B_{g,\alpha}^{\prime\prime}\frac{L^{1/(1+2\alpha)}}{M^{\alpha/(1+3\alpha)}},$
where $A_{g,\alpha}^{\prime},B_{g,\alpha}^{\prime\prime}$ only depends on the properties of $g$ and $\alpha$.
For the remained case, $H<C_{1}C_{2}nq^{2}$, similar to the proof of Theorem 3.3 we can simply choose TF with all 00 parameters. Then the approximation error can be trivially bounded by:
$\left|\!\left|\!\left|\textnormal{{GIH}}_{n}-{\textnormal{{TF}}}\right|\!\right|\!\right|_{L,2}=\left|\!\left|\!\left|\textnormal{{GIH}}_{n}-0\right|\!\right|\!\right|_{L,2}\leq 1.$
Then, the two cases can be unified by:
$\left|\!\left|\!\left|\textnormal{{GIH}}_{n}-{\textnormal{{TF}}}\right|\!\right|\!\right|_{L,2}\leq$$\begin{cases}A_{g,\alpha}^{\prime}\left(\frac{C_{1}nq^{2}}{H}\right)^{q}+B_{g,\alpha}^{\prime\prime}\frac{L^{1/(1+2\alpha)}}{M^{\alpha/(1+3\alpha)}},\quad H\geq C_{1}C_{2}nq^{2}\\ 1,\quad\text{otherwise}\end{cases}$
$\leq$ $\max\{A_{g,\alpha}^{\prime},1\}\left(\frac{C_{1}\max\{C_{2},1\}nq^{2}}{H}\right)^{q}+B_{g,\alpha}^{\prime\prime}\frac{L^{1/(1+2\alpha)}}{M^{\alpha/(1+3\alpha)}}.$
Finally, we can choose $A_{g,\alpha}^{\prime\prime}=\max\{A_{g,\alpha}^{\prime},1\}$ and $C_{n,q}=C_{1}\max\{C_{2},1\}nq^{2}$. Then
$\left|\!\left|\!\left|\textnormal{{GIH}}_{n}-{\textnormal{{TF}}}\right|\!\right|\!\right|_{L,2}\leq A_{g,\alpha}^{\prime\prime}\left(\frac{C_{n,q}}{H}\right)^{q}+B_{g,\alpha}^{\prime\prime}\frac{L^{1/(1+2\alpha)}}{M^{\alpha/(1+3\alpha)}},$
where $A_{g,\alpha}^{\prime\prime}$ and $B_{g,\alpha}^{\prime\prime}$ only depend on $g$ and $\alpha$, thus only depending on the properties of $g$. Moreover, $C_{n,q}=\mathcal{O}(nq^{2})$. ∎
Appendix D Proofs in Section 4
D.1 Optimization Dynamics in Training Stage I